OpenStudy (anonymous):

Help me 4/6-x + x^2/6-x = 2 It already has like denominator I think it is X^2+4/6-x=2

OpenStudy (anonymous):

Yes, it has the like denominator 6-x Therefore, it'll be (4+x^2)/(6-x), or to have it in decreasing power: (x^2+4)/(6-x)=2 Are you just simplifying, or are you solving for x?

OpenStudy (anonymous):

Solving for x

OpenStudy (anonymous):

Ok, so because you have a power to the second (x^2), you're going to get two values for x. 1) get 6-x out of the denominator - so you would multiply both sides by 6-x, resulting in (4+x^2)=2(6-x). You can simplify this so that you have x^2+4=12-2x. 2) After this, you want to set the entire equation equal to zero, so the 12-2x will be moved to the left side via subtraction from the right side 3) With the entire equation set equal to zero, you do a little factoring, which will then give you your two x values in linear form (x-a)(x-b), where a and b are integers. NOTE, in the linear format (x-a) gives an +a value, while (x+a) gives an -a value

OpenStudy (anonymous):

I got x=-2,4 Thanks you